# Conditioning Diffusions Using Malliavin Calculus

### v1

@@ D.2 Equivalence to KL-Loss
$=$ $~{}\mathbb{E}[\|u(X_{t})-\nabla\log p_{t}(Y=y\mid X_{t})\|^{2}\mid Y=y]+\mathbb{E}[\|\mathcal{S}_{t}-\mathbb{E}[\mathcal{S}_{t}\mid X_{t},Y]\|^{2}\mid Y=y]$
@@ D.4 Gaussian Approximations
We will prove more general statements. Indeed, Lemma 3.3 still holds when we replace $X_{T}$ by a general observation $Y\sim q(X_{T},\cdot)$. To that end we first prove that the score is a martingale, when conditioned on the observation $Y$:
Lemma D.5.
Let $s\geqslant t$, then we have that
$\nabla_{X_{t}}\log p(Y\mid X_{t})=\mathbb{E}[\nabla_{X_{t}}\log p_{t}(X_{s}\mid X_{t})\mid Y,X_{t}].$ (50)
Proof.
We have
$\nabla_{x_{t}}\log p(Y\mid X_{t})$
$\qquad=\frac{1}{p(Y\mid X_{t})}\nabla_{x_{t}}\int q(Y\mid X_{T}=x_{T})p(X_{T}=x_{T}\mid X_{t})\mathrm{d}x_{T}$
$\qquad=\frac{1}{p(Y\mid X_{t})}\nabla_{x_{t}}\int q(Y\mid X_{T}=x_{T})p(X_{T}=x_{T}\mid X_{s}=x_{s})p(X_{s}=x_{s}\mid X_{t})\mathrm{d}x_{s}\mathrm{d}x_{T}$
$\qquad=\frac{1}{p(Y\mid X_{t})}\nabla_{x_{t}}\int q(Y\mid X_{T}=x_{T})p(X_{T}=x_{T}\mid X_{s}=x_{s})p(X_{s}=x_{s}\mid X_{t})\nabla_{X_{t}}\log p(X_{s}=x_{s}\mid X_{t})\mathrm{d}x_{s}\mathrm{d}x_{T}$
$\qquad=\mathbb{E}[\nabla_{X_{t}}\log p(X_{s}\mid X_{t})\mid X_{t},Y].$
Now we use the above to show the equivalence of the two losses:
@@
$\mathbb{E}[\|u_{t}(X_{t},Y)-\nabla_{X_{t}}\log p(Y\mid X_{t})\|^{2}]=C+\mathbb{E}[\|u_{t}(X_{t},Y)-\nabla_{X_{t}}\log p(X_{s}\mid X_{t})\|^{2}].$ (51)
@@
$\mathbb{E}[\|u_{t}(X_{t},Y)-\nabla_{X_{t}}\log p(Y\mid X_{t})\|^{2}]=\mathbb{E}[\|u_{t}(X_{t},Y)-\mathbb{E}[\nabla_{X_{t}}\log p(X_{s}\mid X_{t})\mid X_{t},Y]\|^{2}]$
@@
$~{}~{}=\mathbb{E}[\|u_{t}(X_{t},Y)-\mathbb{E}[\nabla_{X_{t}}\log p(X_{s}\mid X_{t})\mid X_{t},Y]\|^{2}]+\mathbb{E}[\|\nabla_{X_{t}}\log p(X_{s}\mid X_{t})-\mathbb{E}[\nabla_{X_{t}}\log p(X_{s}\mid X_{t})\mid X_{t},Y]\|^{2}]$
$~{}~{}=\mathbb{E}[\|u_{t}(X_{t},Y)-\mathbb{E}[\nabla_{X_{t}}\log p(X_{s}\mid X_{t})\mid X_{t},Y]\|^{2}]+\mathbb{E}[\|\nabla_{X_{t}}\log p(X_{s}\mid X_{t})\|^{2}]$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\mathbb{E}[\|[\nabla_{X_{t}}\log p(X_{s}\mid X_{t})\mid X_{t},Y]\|^{2}].$

### v2

@@ D.2 Equivalence to KL-Loss
$=$ $~{}\mathbb{E}[\|u(X_{t})-\nabla\log p(Y=y\mid X_{t})\|^{2}\mid Y=y]+\mathbb{E}[\|\mathcal{S}_{t}-\mathbb{E}[\mathcal{S}_{t}\mid X_{t},Y]\|^{2}\mid Y=y]$
@@ D.4 Gaussian Approximations
Now we use Lemma D.2 to show the equivalence of the two losses:
@@
$\mathbb{E}[\|u_{t}(X_{t},Y)-\nabla_{X_{t}}\log p(Y\mid X_{t})\|^{2}]=C+\mathbb{E}[\|u_{t}(X_{t},Y)-\nabla_{X_{t}}\log p_{s\mid t}(X_{s}\mid X_{t})\|^{2}].$ (51)
@@
$\mathbb{E}[\|u_{t}(X_{t},Y)-\nabla_{X_{t}}\log p(Y\mid X_{t})\|^{2}]=\mathbb{E}[\|u_{t}(X_{t},Y)-\mathbb{E}[\nabla_{X_{t}}\log p_{s\mid t}(X_{s}\mid X_{t})\mid X_{t},Y]\|^{2}]$
@@
$~{}~{}=\mathbb{E}[\|u_{t}(X_{t},Y)-\mathbb{E}[\nabla_{X_{t}}\log p_{s\mid t}(X_{s}\mid X_{t})\mid X_{t},Y]\|^{2}]+\mathbb{E}[\|\nabla_{X_{t}}\log p_{s\mid t}(X_{s}\mid X_{t})-\mathbb{E}[\nabla_{X_{t}}\log p_{s\mid t}(X_{s}\mid X_{t})\mid X_{t},Y]\|^{2}]$
$~{}~{}=\mathbb{E}[\|u_{t}(X_{t},Y)-\mathbb{E}[\nabla_{X_{t}}\log p_{s\mid t}(X_{s}\mid X_{t})\mid X_{t},Y]\|^{2}]+\mathbb{E}[\|\nabla_{X_{t}}\log p_{s\mid t}(X_{s}\mid X_{t})\|^{2}]$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}-\mathbb{E}[\|[\nabla_{X_{t}}\log p_{s\mid t}(X_{s}\mid X_{t})\mid X_{t},Y]\|^{2}].$
@@ Appendix E Gaussian Analysis of the Variance
Appendix F Related Work
Numerous algorithms exist for bridge simulation. For example, Markov chain Monte Carlo-based methods (Delyon & Hu, 2006; Bayer & Schoenmakers, 2013; Bladt & Sørensen, 2014; Schauer et al., 2017) and more recently, machine learning-based works (Heng et al., 2022; Du et al., 2024; Baker et al., 2025; Zhao et al., 2024).
In the context of diffusion models, the most relevant approaches are Zhang et al. (2023); Denker et al. (2024). These methods rely on a fundamental equality in their derivation:
$\nabla_{X_{t}}\log p(X_{T}|X_{t})=\nabla_{X_{t}}\log p(X_{t}|X_{T})+\nabla_{X_{t}}\log p(X_{t})\approx\nabla_{X_{t}}\log p(X_{t}|X_{T})+s_{\theta}(X_{t}).$
This formulation allows the first term on the right to be computed analytically, while the second term relies on the pre-trained score network from the diffusion process. Although this approach offers computational efficiency, it assumes that the score network accurately approximates $\nabla\log p_{t}(X_{t})$ – the gradient of the distribution from which $X_{t}$ is sampled. This assumption frequently fails in practice and may become increasingly invalid after network fine-tuning or other modifications. Our approach avoids this limitation by directly learning optimal control for the current drift estimate $s_{\theta}(X_{t})$, regardless of its origin.
Guidance methods share similar objectives of introducing post-hoc control (Chung et al., 2022; Song et al., 2023; Boys et al., 2023). However, these techniques rely on cheap approximations of the optimal control $\nabla\log p_{t}(Y|X_{t})$, producing samples that deviate from the true posterior or conditional distribution. Instead, they generate samples resembling high-likelihood draws from the prior distribution. While adequate for many applications, these methods cannot provide authentic conditional samples from the Bayesian posterior when such precision is required.
There are also two concurrent works on Malliavin calculus and diffusion models. However, they differ in their goals, techniques and results. In Mirafzali et al. (2025) the authors establish a different Malliavin based score formula. In contrast to Lemma B.1, this construction relies on second-order Malliavin derivatives (Mirafzali et al., 2025, Theorem 3.9) and the inverse of the Malliavin matrix. In Greco (2025) the author uses Malliavin calculus to study diffusion models in infinite dimensions. This can be seen as giving justification for Tweedie’s formula (see 2.5) in the infinite dimensional case, extending the results from Pidstrigach et al. (2024).